Assume a rectangular waveguide with dimensions a=2 cm and b=1cm. The
waveguide is filled with air and the operating frequency is 50MHz. Find the
propagation constant, guide wavelength and cut off frequency for TE10,
TM11 and TE20 modes.
1
Expert's answer
2021-08-03T04:06:01-0400

a)Propagation constant is given by the formula below

"\gamma=\sqrt{({\dfrac{m\pi}{a}})^2+({\dfrac{n\pi}{b}})^2-\omega^2\mu\epsilon}"

a=0.02m,b=0.01m

For TE10 mode, m=1,n=0

f=50MHz

"\gamma=\sqrt{({\dfrac{1\pi}{0.02}})^2+({\dfrac{0\pi}{0.01}})^2-\dfrac{(2\pi\times{50\times{10^6}})^2}{(3\times{10^8})^2}}\\ \gamma=157.1" For TM11 mode

m=1,n=1

"\gamma=\sqrt{({\dfrac{1\pi}{0.02}})^2+({\dfrac{1\pi}{0.01}})^2-\dfrac{(2\pi\times{50\times{10^6}})^2}{(3\times{10^8})^2}}\\ \gamma=351.2" For TE20 mode, m=2, n=0

"\gamma=\sqrt{({\dfrac{2\pi}{0.02}})^2+({\dfrac{0\pi}{0.01}})^2-\dfrac{(2\pi\times{50\times{10^6}})^2}{(3\times{10^8})^2}}\\ \gamma=314.2" The guide wavelength cannot be calculated without cutoff frequency.

"f_c=\dfrac{u'}{2}\sqrt{(\dfrac{m}{a})^2+(\dfrac{n}{b})^2}"

For TE10 mode,m=1,n=0

"f_c=\dfrac{3\times{10^8}}{2}\sqrt{(\dfrac{1}{0.02})^2+(\dfrac{0}{0.01})^2}=7.5GHz"

for TM11 mode,m=n=1

"f_c=\dfrac{3\times{10^8}}{2}\sqrt{(\dfrac{1}{0.02})^2+(\dfrac{1}{0.01})^2}=16.8GHz"for TE20 mode, m=2, n=0

"f_c=\dfrac{3\times{10^8}}{2}\sqrt{(\dfrac{2}{0.02})^2+(\dfrac{0}{0.01})^2}=15MHz"

guide wave is given by

"\lambda_g=\dfrac{c}{\sqrt{f^2-f_c^2}}"

Guide wavelength exist for only TE20 mode

"\lambda_g=\dfrac{3\times10^8}{\sqrt{(50\times{10^6})^2-(15\times{10^6})^2}}=6.29m"